### v1

# Tuning Large Multimodal Models for Videos using Reinforcement Learning from AI Feedback

Abstract
Recent advancements in large language models have influenced the development of video large multimodal models (VLMMs). The previous approaches for VLMMs involved Supervised Fine-Tuning (SFT) with instruction-tuned datasets, integrating LLM with visual encoders, and adding additional learnable modules. Video and text multimodal alignment remains challenging, primarily due to the deficient volume and quality of multimodal instruction-tune data compared to text-only data. We present a novel alignment strategy that employs multimodal AI system to oversee itself called Reinforcement Learning from AI Feedback (RLAIF), providing self-preference feedback to refine itself and facilitating the alignment of video and text modalities. In specific, we propose context-aware reward modeling by providing detailed video descriptions as context during the generation of preference feedback in order to enrich the understanding of video content. Demonstrating enhanced performance across diverse video benchmarks, our multimodal RLAIF approach, VLM-RLAIF, outperforms existing approaches, including the SFT model. We commit to open-sourcing our code, models, and datasets to foster further research in this area. https://github.com/yonseivnl/vlm-rlaif
Figure 1: Qualitative comparison of VLMMs on various video benchmarks. The video question answering task is marked in purple, video-based generative task in blue, the text-to-video retrieval task in violet and the action recognition task in orange color. VLM-RLAIF achieves superior performances on a broad range of video benchmarks compared to previous approaches, including VLM-SFT.
Figure 2: Illustration of the proposed VLM-RLAIF. LLM tuned with video-text instruction-following data, i.e., multimodal SFT model, often produces responses that are not temporally and visually grounded to the video input, as depicted in red color. We propose a method that involves using the VLMM to supervise itself by providing self-preference feedback of generated responses, refining itself and facilitating the alignment of video and text modalities.
1 Introduction
Recent progress in large language models (LLMs) has advanced the field of video large multimodal models (VLMMs) Li et al. (2023a); Muhammad Maaz and Khan (2023); Lin et al. (2023). By extending the logical reasoning and advanced cognitive capabilities of LLMs to the visual domain, these advancements have enabled VLMMs to showcase remarkable proficiency in tasks such as video understanding Li et al. (2023a), video question answering Ko et al. (2023) and instruction-following tasks Muhammad Maaz and Khan (2023); Luo et al. (2023).
These VLMMs achieve this by integrating pre-trained LLMs, such as LLaMA Touvron et al. (2023), with visual encoders and an additional learnable module, i.e., LoRA Hu et al. (2022). Specifically, they undergo a supervised fine-tuning (SFT) stage using multimodal instruction-tune data Luo et al. (2023); Muhammad Maaz and Khan (2023); Li et al. (2023a). This process aims to adapt LLMs to video modality, thus improving their ability to interpret visual content.
However, there is a challenge in achieving seamless multimodal alignment between video and text, primarily due to the deficiency in volume and quality of multimodal instruction-tune data compared to text-only data. Text-only data are typically abundant and diverse, while multimodal data are often limited in both quantity and comprehensiveness Wei et al. (2021); Liu et al. (2023a). This often leads the VLMMs to generate responses that are not properly grounded in the visual context, as illustrated by the SFT model responses in Fig. 2.
To address this, we propose a novel alignment approach for video and text modality that involves using the VLMM to supervise itself by providing preference feedback of generated responses, as shown in Fig. 2. Specifically, we introduce Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022); Lee et al. (2023) for multimodal alignment. Unlike Reinforcement Learning with Human Feedback (RLHF), which has been successful in aligning text-only or image-text based AI agents with human intentions Ouyang et al. (2022); Sun et al. (2023a); Yu et al. (2023), the RLAIF allows for scalable oversight with minimum human intervention. Particularly, we enhance AI’s feedback by proposing a context-aware reward modeling, focusing on improved clarity in video content (Sec. 3.2). Furthermore, to compensate for the limited multimodal instruction-tune data for training the SFT, we augment it with a human-labeled video question answering and an object-centric multimodal instruction-tune dataset. Further, to effectively utilize this expanded dataset, we propose a simple curriculum training strategy for enhancing the alignment between video and text modality (Sec. 3.3).
To sum up, we call our proposed framework of training an VLMM with AI feedback as Video Large Multimodal model with RLAIF or VLM-RLAIF for short. Our empirical studies show that our aligned VLM-RLAIF exhibits superior performance compared to state-of-the-art VLMMs across a wide array of video benchmarks, as illustrated in Fig. 1.
We summarize our contributions as follows:
  • We propose a novel alignment method for video large multimodal models (VLMMs), utilizing Reinforcement Learning from AI feedback (RLAIF) to align video-text modalities effectively.
  • We enhance AI’s feedback by proposing context-aware reward modeling, focusing on improved clarity and comprehension in video.
  • We enrich the SFT model’s training by introducing additional instruction-tune data and applying a simple curriculum strategy.
  • We demonstrate the effectiveness of our proposed VLM-RLAIF on various video understanding benchmarks by a noticeable margin.
2 Related Work
Multimodal large model.
In recent research efforts, there has been a significant focus on advancing Language Models (LLMs) to go beyond natural language understanding, extending into the realm of multi-modal comprehension. The goal is to develop LLMs capable of interpret additional modalities, e.g., image Liu et al. (2023a), video Li et al. (2023a); Lin et al. (2023), 3D point-cloud Guo et al. (2023), etc.
To provide additional modality into LLMs, most of work utilize pre-trained encoder, such as CLIP Radford et al. (2021), Q-former Li et al. (2022) or ImageBind Girdhar et al. (2023), to extract each modality feature from data. These features then projected to the token embedding space of the language model. Following this, the models are subject to supervised fine-tuning (SFT) using synthetically generated, modality-specific instruction-following datasets. These approaches, exemplified in works such as LLaVA Liu et al. (2023a), Video-LLaVA Lin et al. (2023) or Point-LLM Guo et al. (2023), facilitates the development of models proficient in engaging in conversations grounded in additional modality.
Figure 3: Detailed illustration of the three stages for context-aware reward modeling. Dotted box in each stage represents distinct model input: the first is for generating responses using the SFT model, the second for the judge model to evaluate and choose the superior response between options A and B, and the third for training the RM. Each input includes a task-specific prompt, denoted by the yellow box, tailored to guide the model’s operation within its respective function (more details are in Appendix Sec. C). The operational flow of each stage is as follows: (1) The SFT model produces two candidate responses from the provided video and question. (2) With the video, question and responses at hand, the SFT model utilize context information and guiding prompt to evaluate the responses. (3) The RM is trained using the preference pairs generated in the previous step as indicated in orange box.
Figure 4: An example of context-aware preference selection. We demonstrate our model’s ability to generate preference feedback, i.e., preferred choice and explanation marked in orange box, on given instruction input, two responses and context. Red color denotes an incorrect response, while blue color signifies a correctly grounded response with respect to the visual input. The rationale behind selecting ‘Output B’ as the preferred choice is indicated in green.
Reinforcement learning from feedback.
To operate the model safely and in accordance with human intentions, Reinforcement Learning from Human Feedback (RLHF) has been proposed as a viable solution Ouyang et al. (2022); Sun et al. (2023a). By collecting the preferences from human evaluators, it usually trains the reward model that gives a high reward to the preferred output of the model. However, significant challenge in this process is the annotation cost associated with selecting preference. To mitigate this issue, the concept of Reinforcement Learning with AI-inferred Feedback (RLAIF) was introduced Bai et al. (2022); Lee et al. (2023); Sun et al. (2023b). RLAIF capitalizes on the inherent ability of Large Language Models (LLMs) to rate or evaluate information, allowing the LLM itself to assign preferences.
3 VLM-RLAIF Framework
3.1 Overview
Contrasting to the limited scalability of human feedback in RLHF, we leverage AI’s feedback for aligning multimodality between video and text, reducing the reliance on exhaustive human-annotated preferences Ouyang et al. (2022); Sun et al. (2023a). The training procedure of VLM-RLAIF can be summarized into three stages as follows:
Supervised fine-tuning (SFT).
We fine-tune an LLM, i.e., Vicuna, using supervised learning on synthetically generated video-text instruction-tune data Muhammad Maaz and Khan (2023). This involves integrating a vision encoder with two linear layers and an additional learnable parameters using LoRA Hu et al. (2022), into the training process. Such fine-tuning enhances the model’s capability to follow instructions. We refer to this model as the Video Large Multimodal model with SFT or VLM-SFT for short.
Reward modeling with AI feedback.
A key aspect of the RLAIF involves leveraging a pre-trained AI model to generate human-like preferences between different responses generated from the same input Bai et al. (2022); Sun et al. (2023b); Lee et al. (2023). To accomplish this, we employ the VLM-SFT as a judge for assessing preferences. Once preferences are annotated, we proceed to train a reward model (RM) based on these annotated labels using a cross-entropy loss, following the Bradley-Terry model for estimating score functions from pairwise preferences Ouyang et al. (2022); Sun et al. (2023a). We illustrate the training procedure for collecting preferences and training the reward model in the Fig. 3. The RM give higher score reward to the better response and lower score reward to the less appropriate one in a pair of responses, thus providing guidance to a policy model during the reinforcement learning (RL) process.
Reinforcement learning.
During the RL phase, we fine-tune a supervised policy model, initialized from the VLM-SFT, aiming to optimize the scalar reward output from the trained RM. We use the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) for this purpose, following prior research Ouyang et al. (2022); Sun et al. (2023b, a). This process finely tunes the model to respond optimally to the rewards, thereby enhancing its decision-making abilities in accordance with the learned preferences.
3.2 Context-Aware Reward Modeling
For effective utilization of VLM-SFT in preference selection, a nuanced understanding of video content is essential, particularly in scenarios involving complex scenes with elements subject to varied interpretations. Acknowledging this intricacy, it becomes apparent that exclusive dependence on VLM-SFT for direct video analysis may fall short in achieving thorough comprehension.
Context-aware preference selection.
In response, we enhance the VLM-SFT by integrating detailed video descriptions, referred to as context, into the preference selection workflow, thereby imparting additional contextual clarity, as illustrated in Figures 3-(2) and 4. This process entails breaking down the video into smaller clips, followed by generating descriptive caption for each segment to obtain fine-grained information. Subsequently, these individual captions are concatenated into an exhaustive narrative of the video. This detailed narrative is then provided to the judge model, i.e., VLM-SFT, to aid in better preference selection. This process not only improves the VLM-SFT’s ability to obtain a comprehensive view of the video content, but also enables it to more identify the most suitable response. Integrating all inputs, i.e., marked in dotted box in Fig. 3-(2), with a specific prompt, i.e., rules for generating preferences as illustrated in Appendix Fig. 9, allows us to collect context-aware synthetic preferences, thus enriching the model’s decision-making process.
Training reward model.
The reward model (RM) is designed to assign higher scores to responses considered better and lower scores to those considered lower in quality. Starting from VLM-SFT with the final linear layer removed, we train a model to take in an input prompt and response, i.e., marked in dotted box of Fig. 3-(3), and output a scalar reward value. Using the preference dataset produced from the VLM-SFT, we train the RM with a cross-entropy loss. Specifically, we use 13B VLM-SFT to train the reward model, as it gives slightly better performance than 7B VLM-SFT (see the quantitative comparison discussed in Appendix Sec. A). Remarkably, after training RLAIF using the RM, the 7B VLM-RLAIF significantly surpass the 13B VLM-SFT, thus validating the effectiveness of our proposed framework in aligning video and text modalities (refer to Sec. 4.2 for more details).
3.3 Two-stage Curriculum Supervised Fine-Tuning
To obtain video conversational capability, we train the LLM with a open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023). Further, to enhance this, we not only expand our training with additional video-text instruction-tune datasets but also introduce a simple curriculum learning strategy. This approach is designed to improve the model’s performance progressively, ensuring a more effective and nuanced understanding of video-text interactions.
Augmenting video instruction-tune dataset.
To improve the video understanding ability, we first augment the open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023) with existing human-annotated video question answering datasets Xiao et al. (2021); Li et al. (2020). Moreover, recognizing the importance of aligning LLMs with the video modality, we focus on obtaining instruction-tune dataset that encompass both visual and temporal intricacies. To achieve this, we leverage Video Localized Narratives Voigtlaender et al. (2023), a dataset that provides detailed descriptions of the appearance and complex actions of all objects in the video. We transform the narrative dataset into object-centric instruction-tune dataset format using ChatGPT similar to prior work Muhammad Maaz and Khan (2023). More details about the generated instruction-tune dataset can be found in the Appendix Sec. B. We train the SFT model using both the open-sourced and our proposed datasets, resulting in improved performance, as discussed in the Sec. 4.3.
Curriculum supervised fine-tuning
Recent VLMM research has advanced instruction-following capabilities through curating multimodal instruction-tune data Muhammad Maaz and Khan (2023); Luo et al. (2023), joint learning across image and video datasets Lin et al. (2023); Li et al. (2023c), and enhanced bridging networks like adapter Liu et al. (2023b). Unlike these approaches, we emphasize a structured learning method for VLMMs, inspired by the human learning process. We advocate for the importance of the training sequence, drawing from curriculum learning principles Chang et al. (2021), and propose a simple two-stage curriculum learning strategy for SFT, aiming to optimize the instruction-following proficiency of VLMMs.
To accomplish this, we divide the SFT dataset based on complexity into two segments: ‘easy’ and ‘hard’. The model is first trained on ‘easy’ tasks to learn basic concepts, then progressed to ‘hard’ tasks for advanced challenges. Given that longer answers often require deeper comprehension of the context and enhanced proficiency in creating syntactically complex sentences Xu et al. (2020); Agrawal and Singh (2023); Ranaldi et al. (2023), we use answer length as our proxy for sample difficulty, meaning that the longer the answer sentence, the more difficult the task is. This proxy also allows for a clear basis for dividing data, offering a concrete measure of difficulty that aligns the model’s learning trajectory with the incremental nature of human learning.
Methods LLM Size Video-based Generative Performance
Correctness Detail Context Temporal Consistency
VideoChat Li et al. (2023a) 7B 2.23 2.50 2.53 1.94 2.24
LLaMA-Adapter Zhang et al. (2023b) 7B 2.03 2.32 2.30 1.98 2.15
VideoLLaMA Zhang et al. (2023a) 7B 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT Muhammad Maaz and Khan (2023) 7B 2.40 2.52 2.62 1.98 2.37
Valley Luo et al. (2023) 7B 2.43 2.13 2.86 2.04 2.45
BT-Adapter Liu et al. (2023b) 7B 2.68 2.69 3.27 2.34 2.46
VTimeLLM Huang et al. (2023) 7B 2.78 3.10 3.40 2.49 2.47
Video-LLaVA Lin et al. (2023) 7B 2.84 2.86 3.44 2.46 2.57
VideoChat2 Li et al. (2024) 7B 3.02 2.88 3.51 2.66 2.81
LLaMA-VID Li et al. (2023c) 7B 2.96 3.00 3.53 2.46 2.51
LLaMA-VID Li et al. (2023c) 13B 3.07 3.05 3.60 2.58 2.63
GPT-4V OpenAI et al. (2023) - 3.85 3.45 3.84 3.63 2.8
VLM-SFT 7B 2.79 2.82 3.37 2.28 2.49
VLM-RLAIF 7B 3.63 3.25 4.00 3.23 3.32
$\Delta$ (RLAIF - SFT) - +0.84 +0.43 +0.63 +0.95 +0.83
Table 1: Qualitative comparison between different VLMMs on video-based generative performance benchmark. We assess various VLMMs on video-based generative benchmark which measures various generative aspects. Our approach, VLM-RLAIF, shows a performance improvement over previous approaches, with the exception of GPT-4V, and demonstrates noticeable enhancements across five criteria when compared to the VLM-SFT. Here, $\Delta$ (RLAIF - SFT) indicates the improvement of RLAIF model over SFT model. $\dagger$} denotes reproduced results using the author’s implementation.
4 Experiments
4.1 Experimental Setup
Model settings.
We initiate training of the VLM-SFT, based on the pre-trained LLaVA image-text large multimodal models Liu et al. (2023a), leveraging a diverse video-text instruction-tune dataset Muhammad Maaz and Khan (2023); Xiao et al. (2021); Li et al. (2020); Voigtlaender et al. (2023). This training incorporates a curriculum approach as outlined in Sec. 3.3. Upon establishing the VLM-SFT, we proceed to train the reward model (RM) using the VLM-SFT for both its foundation and the generation of preference data. Subsequently, we implement reinforcement learning with the trained RM, leading to the development of the VLM-RLAIF. In the reinforcement learning phase, the policy model is initialized from the VLM-SFT, while the value model employs the RM.
Dataset details.
For the SFT dataset, we utilize the open-sourced video-text instruction-tune dataset (80k) Muhammad Maaz and Khan (2023); Li et al. (2023a) and video question answering datasets (67k) Xiao et al. (2021); Li et al. (2020). Additionally, we generate object-centric narrative video-text instruction-tune dataset (180k) for training the VLM-SFT, as mentioned in Sec. 3.3. For two-stage curriculum learning, we divide the instruction-tune data for SFT into two groups based on the difficulty; easy (214k) and hard (113k) data. To train the reward model (RM), we first generate responses from existing instruction-tune data Muhammad Maaz and Khan (2023) and generate preferences using them (40k). Then, we once again utilize the existing instruction-tune dataset (100k) Muhammad Maaz and Khan (2023) for reinforcement learning with the trained RM for initialization.
Evaluation details.
We evaluate our VLM-RLAIF on various video benchmarks; zero-shot video questions and video-based generative benchmarks following the prior work Muhammad Maaz and Khan (2023); Lin et al. (2023); Li et al. (2023c) and text-to-video-retrieval and action recognition as proposed in prior work Li et al. (2023b).
Training details.
[…]
Methods LLM Size MSVD-QA MSRVTT-QA ActivityNet-QA
Acc. Score Acc. Score Acc. Score
FrozenBiLM Yang et al. (2022) 1B 32.2 - 16.8 - 24.7 -
VideoChat Li et al. (2023a) 7B 56.3 2.8 45.0 2.5 26.5 2.2
LLaMA-Adapter Zhang et al. (2023b) 7B 54.9 3.1 43.8 2.7 34.2 2.7
VideoLLaMA Zhang et al. (2023a) 7B 51.6 2.5 29.6 1.8 12.4 1.1
Video-ChatGPT Muhammad Maaz and Khan (2023) 7B 64.9 3.3 49.3 2.9 35.2 2.7
Valley Luo et al. (2023) 7B 60.5 3.3 51.1 2.9 45.1 3.2
BT-Adapter Liu et al. (2023b) 7B 67.5 3.7 57.0 3.2 45.7 3.2
Video-LLaVA Lin et al. (2023) 7B 70.7 3.9 59.2 3.5 45.3 3.3
VideoChat2 Li et al. (2024) 7B 70.0 3.9 54.1 3.3 49.1 3.3
LLaMA-VID Li et al. (2023c) 7B 69.7 3.7 57.7 3.2 47.4 3.3
LLaMA-VID Li et al. (2023c) 13B 70.0 3.7 58.9 3.2 47.5 3.3
VLM-SFT 7B 67.2 3.6 52.4 3.0 44.1 3.2
VLM-RLAIF 7B 76.4 4.0 63.0 3.4 57.3 3.5
$\Delta$ (RLAIF - SFT) - +9.2% +0.4 +10.6% +0.4 +13.2% +0.3
Table 2: Qualitative comparison between different VLMMs on zero-shot video question answering benchmark. Our VLM-RLAIF outperforms previous work across three video-question answering benchmarks.
Methods LLM Size T2V Retrieval Action Recognition
MSVD MSRVTT UCF101 HMDB51
R@1 R@5 R@1 R@5 Top-1 Top-5 Top-1 Top-5
Video-ChatGPT Muhammad Maaz and Khan (2023) 7B 26.03 51.25 14.60 33.80 51.49 79.25 37.10 63.97
Video-LLaVA Lin et al. (2023) 7B 29.34 55.35 18.70 38.60 52.33 80.86 36.64 64.03
LLaMA-VID Li et al. (2023c) 7B 27.28 53.40 17.00 35.10 56.58 82.79 38.85 65.27
VLM-SFT 7B 26.65 54.27 13.10 30.50 53.03 80.34 38.58 62.37
VLM-RLAIF 7B 36.03 63.40 21.00 40.70 62.83 85.86 44.75 68.37
$\Delta$ (RLAIF - SFT) - +9.38 +9.13 +7.90 +10.2 +9.80 +5.52 +8.11 +6.00
Table 3: Qualitative comparison between different VLMMs on zero-shot text-to-video (T2V) retrieval and action recognition. Following prior work Li et al. (2023b), we assess our proposed VLM-RLAIF on zero-shot T2V and action recognition for comprehensive evaluation of the VLMMs. $\dagger$} denotes reproduced results using the author’s implementation.
Figure 5: Effect of various preference data size on video-based generative benchmark. VLM-RLAIF performance increase with the number of collected preference data.
4.2 Quantitative Analysis
We quantitatively evaluate our proposed VLM-RLAIF on various video benchmarks following prior work: video-based generative benchmark, zero-shot video question answering Muhammad Maaz and Khan (2023); Lin et al. (2023); Li et al. (2023c), text-to-video retrieval and action recognition Li et al. (2023b).
Video-based generative performance.
We evaluate VLMMs on the video-based generative performance benchmark Muhammad Maaz and Khan (2023) that measures five criteria: correctness, detail, context, temporal understanding, and consistency of generated text. In specific, these assess the relevance of the model’s output to the video content, its capacity to capture essential details and contextual information, its understanding of temporal sequences, and the consistency in responding to varied yet related queries. As indicated in Tab. 1, the VLM-RLAIF performs on par with GPT-4V OpenAI et al. (2023) and outperforms previous approaches and the VLM-SFT.
Zero-shot video question answering.
To assess the general capabilities of various video-text Large Language Models (LLMs), we conduct a quantitative evaluation of their video question answering (VideoQA) abilities using three datasets: MSVD-QA, MSRVTT-QA Xu et al. (2017), and ActivityNet-QA Yu et al. (2019). We adopted the evaluation pipeline from prior work Muhammad Maaz and Khan (2023), which utilizes ChatGPT-assisted evaluation, to gauge the zero-shot VideoQA performance. The results, as shown in Table 2, indicate that the VLM-RLAIF significantly outperforms previous models. Notably, VLM-RLAIF exceeds VLM-SFT by 9.2%, 10.6%, and 13.2% in accuracy and by 0.4, 0.4, and 0.3 in score across all datasets.
Zero-shot text-to-video retrieval.
For this evaluation, we follow the procedure proposed in prior work Li et al. (2023b), which compare CLIP score between generated description and ground-truth caption. Table 3 illustrates the summarized performance comparison to various VLMMs. In two datasets, i.e., MSVD and MSRVTT, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
Zero-shot action recognition.
Following the VLMMs evaluation procedure proposed in prior work Li et al. (2023b), we conduct zero-shot action recognition task using two benchmark datasets, e.g., UCF101 and HMDB51. Table 3 illustrates the summarized performance comparison to various VLMMs. In two datasets, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
SFT datasets Curr. learning Video-based Generative Performance
[A] [B] [C] Corr. Det. Cont. Temp. Cons.
2.32 2.53 3.03 2.16 2.23
2.43 2.56 3.09 2.19 2.19
2.79 2.82 3.37 2.28 2.49
Table 4: In-depth analysis for the VLM-SFT training procedure. ‘[A]’ indicate the multimodal instruction-tune dataset proposed in Muhammad Maaz and Khan (2023); Li et al. (2023a). ‘[B]’ represents the use of a human-labeled video question answering dataset Xiao et al. (2021); Li et al. (2020), while ‘[C]’ refers to the use of an object-centric video narrative instruction-tune dataset (Appendix Sec. B). ‘Curr. learning’ indicates the application of curriculum learning as detailed in Sec. 3.3.
RLAIF Context #\#Clips Video-based Generative Performance
Info. Corr. Det. Cont. Temp. Cons.
- 2.79 2.82 3.37 2.28 2.49
- 3.26 3.11 3.74 2.78 3.14
1 3.44 3.20 3.89 2.97 3.36
3 3.63 3.25 4.00 3.23 3.32
Table 5: Effect of context information on video-based generative performance benchmark Muhammad Maaz and Khan (2023). We investigate the efficacy of using context information for reward modeling (Sec. 3.2). ‘Context Info.’ indicates the use of context in preference selection. ‘#\# Clips’ refers to the the number of times we divide the video into segments to generate the context information.
4.3 Detailed Analysis
For a detailed analysis, we leverage the video-based generative benchmark Muhammad Maaz and Khan (2023) specifically, as it is well suited to assess the wide-ranging capabilities of VLMMs.
In-depth analysis of SFT training.
We first empirically support the effectiveness of augmenting the SFT dataset with additional instruction-following dataset (Sec. 3.3). The first and second rows of Tab. 4 illustrate the benefits of incorporating this additional dataset in improving video-based generative performance. On top of that, the application of curriculum learning significantly boosts performance, demonstrating the efficacy of employing a curriculum learning strategy during the SFT process, as illustrated in the third row of Tab. 4.
Effect of various preference data size.
Our framework’s strength lies in generating synthetic preference feedback in large quantities. To evaluate this capability, we escalated the amount of preference data from 10k – the same quantity utilized in the multimodal image-text RLHF framework Sun et al. (2023a) – to 40k. Figure 5 shows a steady improvement in performance with the increase in data size.
Context-aware preference labeling.
Table 5 presents a study on the impact of context information on the performance of the video-based generative benchmark discussed in Sec. 3.2. This reveals that the use of context information during preference selection improves model performance, as demonstrated in the second and third rows in Table 5. Moreover, segmenting the video into multiple clips, specifically three, and using their aggregated caption as context information further enhances performance, as indicated in the third and fourth rows. We conjecture that segmenting video into clips and providing detailed captions for each clip enables the model to capture more detailed nuances, thereby enriching the contextual clarity for reward modeling and boosting the VLM-RLAIF’s performance.
Figure 6: Samples of VLM-RLAIF and VLM-SFT in video understanding. Two samples demonstrate superior capability of the VLM-RLAIF. Red color indicates visually un-grounded expressions considering video content and blue color well grounded to the visual content.
4.4 Qualitative Analysis
VLM-SFTvs. VLM-RLAIF.
Figure 6 qualitatively compares the performance of VLM-SFT and VLM-RLAIF across two video samples, highlighting their multimodal understanding capabilities. VLM-RLAIF consistently yields more accurate answers than VLM-SFT, as seen in the detailed recognition of an athlete’s attire in the first example, marked in red and blue. The second example further affirms VLM-RLAIF’s superiority in generating precisely grounded responses to the visual input, where VLM-SFT falls short. More examples are in the Appendix Fig. 11.
5 Conclusion
We introduce a novel alignment strategy for VLMMs, termed VLM-RLAIF, utilizing reinforcement learning from AI feedback. To enhance multimodal alignment, we also present a context-aware reward modeling technique, enabling AI to generate feedback for self-improvement. Additionally, we expand the instruction-tune dataset for SFT and adopt a straightforward curriculum training approach. The VLM-RLAIF significantly surpasses previous models across multiple multimodal video-text understanding benchmarks, including video-based generative tasks, zero-shot VideoQA, text-to-video retrieval, and action recognition.
Appendix A Quantitative Comparison between Different SFT Model Sizes
We assess the performance of the VLM-SFT using varying LLM sizes, specifically 7B and 13B, across video question answering (VideoQA) and video-based generative benchmarks, detailed in Table 6. The results indicate similar outcomes in the zero-shot VideoQA tasks across both model sizes. However, in video-based generative tasks, a clear trend emerges: larger LLM sizes lead to improved performance. This suggests that the increased model capacity inherent in larger LLMs enhances the model’s ability to capture and generate more complex and nuanced video content representations. Thus, we adopt the 13B model for the reward model (Sec. 3.2), with empirical evidence provided in Tables 7 and 8.
Appendix B Details About Object-Centric Instruction-Tune Data
To enhance the alignment of Large Language Models (LLMs) with video content, we prioritize acquiring more video-text instruction-tune dataset that captures both the visual and temporal complexities of videos. For this, we leverage the Video Localized Narratives dataset Voigtlaender et al. (2023), which contains a comprehensive description of each object’s appearance and action, along with their tracked coordinates within videos. To utilize the best of the rich semantics contained in narrative dataset, we feed all ground-truth captions to the model, i.e., ChatGPT, and prompt it to generate question and answer pairs, as shown in the template in Fig. 7-(a). We aim to transform all object’s descriptions into instruction-tune data format, i.e., question and answer pair, which demands an understanding of the visual specifics of each object and its surroundings, ensuring the questions and answers are anchored in the video. An example of the generated instruction-tune data is depicted in Fig.7-(b), showcasing the approach’s effectiveness in creating contextually rich instructional content.
Appendix C Input Prompts for Preference Modeling
We leverage three types of input prompts for context-aware preference modeling, as discussed in Sec. 3.2. Figures 89 and 10 show the detailed prompt example for each stage, respectively.
Appendix D Various LLM Size for Reward Model
In Tables 7 and 8, we assess the performance of VLM-RLAIF using varying reward model(RM) sizes, each initialized from VLM-RLAIF-7B and VLM-RLAIF-13B. In both settings, we use 7B VLM-SFT to initialize the policy model. For both sizes of RM, our RLAIF method showed a performance gain with a large margin from VLM-SFT across all evaluation methods. Specifically, RLAIF with 7B RM acquired 5-12% improvement on zero-shot video question answering. Scaling up the RM from 7B to 13B result in more improved performance, except for text-to-video retrieval task R@5 metric on MSR-VTT.
Figure 7: Designed input prompt for object-centric instruction-tune dataset generation from video narrative dataset and the data sample. In (a), the prompt we used to generate instruciton-following dataset is displayed in a white box. It includes the system prompt, a task definition, guidelines, principles, and an example set of input and desired output. Using this prompt with ChatGPT-3.5-turbo, we create an additional instruction following datasets. The resulting example is visualized in (b). The question covers details of the video scenes, such as the number and appearance of parrots and the action of each parrot.
Figure 8: Designed input prompt fed to an VLM-SFT to generate responses.
Figure 9: Designed input prompt fed to an VLM-SFT to generate AI’s preference feedback.
Figure 10: Designed input prompt fed to a reward model to generate scalar reward score.
Figure 11: More examples of the comparative results between VLM-SFT and VLM-RLAIF. We present more qualitative examples of questions and answers generated by the VLM-SFT and VLM-RLAIF. The VLM-RLAIF yields responses that are more accurately grounded in the visual input, as highlighted in blue for accurate responses and red for less accurate ones from VLM-SFT.

### v3

@@
 Recent advancements in large language models have influenced the development of video large multimodal models (VLMMs).
-The previous approaches for VLMMs involved Supervised Fine-Tuning (SFT) with instruction-tuned datasets, integrating LLM with visual encoders, and adding additional learnable modules.
-Video and text multimodal alignment remains challenging, primarily due to the deficient volume and quality of multimodal instruction-tune data compared to text-only data.
-We present a novel alignment strategy that employs multimodal AI system to oversee itself called Reinforcement Learning from AI Feedback (RLAIF), providing self-preference feedback to refine itself and facilitating the alignment of video and text modalities.
-In specific, we propose context-aware reward modeling by providing detailed video descriptions as context during the generation of preference feedback in order to enrich the understanding of video content.
-Demonstrating enhanced performance across diverse video benchmarks, our multimodal RLAIF approach, VLM-RLAIF, outperforms existing approaches, including the SFT model.
+Previous approaches for VLMMs involve Supervised Fine-Tuning (SFT) with instruction-tuned datasets, integrating LLM with visual encoders, and additional learnable parameters.
+Here, aligning video with text, and vice versa, remains a challenge, primarily due to the insufficient quality and quantity of multimodal instruction-tune data compared to that of text-only.
+This discrepancy often results in alignments that poorly ground the video content.
+To address this, we present a novel alignment strategy that employs a multimodal AI system equipped with Reinforcement Learning from AI Feedback (RLAIF), providing self-preference feedback to refine itself and facilitating the alignment of video and text modalities.
+Our approach uniquely integrates detailed video descriptions as context into a multimodal AI system during preference feedback generation to enrich the understanding of video content, a process we call context-aware reward modeling.
+Empirical evaluations on various video benchmarks demonstrate that our VLM-RLAIF outperforms existing approaches, including the SFT model.
 We commit to open-sourcing our code, models, and datasets to foster further research in this area.
 https://github.com/yonseivnl/vlm-rlaif
-Figure 1: Qualitative comparison of VLMMs on various video benchmarks.
-The video question answering task is marked in purple, video-based generative task in blue, the text-to-video retrieval task in violet and the action recognition task in orange color.
+Figure 1: Quantitative comparison of VLMMs on various video benchmarks.
+The video question answering (VQA) task is marked in purple, video-based generative task in blue, the text-to-video (T2V) retrieval task in violet and the action recognition (AR) task in orange color.
 VLM-RLAIF achieves superior performances on a broad range of video benchmarks compared to previous approaches, including VLM-SFT.
-Figure 2: Illustration of the proposed VLM-RLAIF.
-LLM tuned with video-text instruction-following data, i.e., multimodal SFT model, often produces responses that are not temporally and visually grounded to the video input, as depicted in red color.
-We propose a method that involves using the VLMM to supervise itself by providing self-preference feedback of generated responses, refining itself and facilitating the alignment of video and text modalities.
+Comprehensive comparisons are provided in Tables 1, 2 and 3.
 1 Introduction
-Recent progress in large language models (LLMs) has advanced the field of video large multimodal models (VLMMs) Li et al.
-(2023a); Muhammad Maaz and Khan (2023); Lin et al.
+Large language models (LLMs) are advancing many language and multimodal AI tasks, including those involved with video large multimodal models (VLMMs) Li et al.
+(2023b); Muhammad Maaz and Khan (2023); Lin et al.
 (2023).
-By extending the logical reasoning and advanced cognitive capabilities of LLMs to the visual domain, these advancements have enabled VLMMs to showcase remarkable proficiency in tasks such as video understanding Li et al.
-(2023a), video question answering Ko et al.
+Extending the logical reasoning and advanced cognitive capabilities of LLMs to the visual domain, VLMMs are now remarkably proficient in tasks such as video understanding Li et al.
+(2023b), video question answering Ko et al.
 (2023) and instruction-following tasks Muhammad Maaz and Khan (2023); Luo et al.
 (2023).
-These VLMMs achieve this by integrating pre-trained LLMs, such as LLaMA Touvron et al.
-(2023), with visual encoders and an additional learnable module, i.e., LoRA Hu et al.
-(2022).
-Specifically, they undergo a supervised fine-tuning (SFT) stage using multimodal instruction-tune data Luo et al.
+These models include publicly available LLMs Touvron et al.
+(2023); Chiang et al.
+(2023); Taori et al.
+(2023) with visual encoders and additional learnable parameters Hu et al.
+(2022); Liu et al.
+(2023b); Li et al.
+(2023a).
+To adapt LLMs to the video modality, thus improving their ability to interpret visual content, they all undergo a supervised fine-tuning (SFT) stage using multimodal instruction-tune data Luo et al.
 (2023); Muhammad Maaz and Khan (2023); Li et al.
-(2023a).
-This process aims to adapt LLMs to video modality, thus improving their ability to interpret visual content.
-However, there is a challenge in achieving seamless multimodal alignment between video and text, primarily due to the deficiency in volume and quality of multimodal instruction-tune data compared to text-only data.
-Text-only data are typically abundant and diverse, while multimodal data are often limited in both quantity and comprehensiveness Wei et al.
+(2023b).
+Figure 2: Illustration of the proposed VLM-RLAIF.
+An LLM tuned with video-text instruction-following data, i.e., multimodal SFT model, often produces responses that are not temporally and visually grounded to the video input, as depicted in red color.
+We propose a method that involves using the VLMM to supervise itself by providing self-preference feedback of generated responses using reward model, refining itself and facilitating the alignment of video and text modalities.
+However, multimodal alignment between video and text faces a significant challenge of deficiency in volume and quality of multimodal instruction-tune data compared to text-only data; text-only data are typically abundant and diverse, while multimodal data are often limited in both quantity and comprehensiveness Wei et al.
 (2021); Liu et al.
 (2023a).
-This often leads the VLMMs to generate responses that are not properly grounded in the visual context, as illustrated by the SFT model responses in Fig.
+This often leads the VLMMs to generate responses that are not properly grounded in the visual content, as illustrated by the SFT model responses in Fig.
 2.
-To address this, we propose a novel alignment approach for video and text modality that involves using the VLMM to supervise itself by providing preference feedback of generated responses, as shown in Fig.
+To address the issue of VLMMs producing inadequately grounded response, we propose a novel method to align video with text that involves using the VLMM to supervise itself by providing preference feedback of generated responses, as shown in Fig.
 2.
-Specifically, we introduce Reinforcement Learning from AI Feedback (RLAIF) Bai et al.
+Specifically, we propose to use Reinforcement Learning from AI Feedback (RLAIF) Bai et al.
 (2022); Lee et al.
 (2023) for multimodal alignment.
-Unlike Reinforcement Learning with Human Feedback (RLHF), which has been successful in aligning text-only or image-text based AI agents with human intentions Ouyang et al.
+Unlike the Reinforcement Learning with Human Feedback (RLHF), which has been successful in aligning text-only or image-text based AI agents with human intentions Ouyang et al.
 (2022); Sun et al.
 (2023a); Yu et al.
-(2023), the RLAIF allows for scalable oversight with minimum human intervention.
-Particularly, we enhance AI’s feedback by proposing a context-aware reward modeling, focusing on improved clarity in video content (Sec.
+(2023), the RLAIF allows for scalable oversight with minimal human intervention.
+In particular, we improve AI feedback by leveraging detailed video descriptions as a context during the generation of preference feedback, focusing on improved clarity in video content, a process we refer to as context-aware reward modeling (Sec.
+3.1).
+In addition, to compensate for the limited multimodal instruction-tune data for training the SFT, we augment it with a human-labeled video question answering and an object-centric multimodal instruction-tune dataset.
+Further, to effectively utilize this expanded dataset, we propose a simple curriculum training strategy for enhancing the alignment between video and text modality (Sec.
 3.2).
-Furthermore, to compensate for the limited multimodal instruction-tune data for training the SFT, we augment it with a human-labeled video question answering and an object-centric multimodal instruction-tune dataset.
-Further, to effectively utilize this expanded dataset, we propose a simple curriculum training strategy for enhancing the alignment between video and text modality (Sec.
-3.3).
-To sum up, we call our proposed framework of training an VLMM with AI feedback as Video Large Multimodal model with RLAIF or VLM-RLAIF for short.
+We call our proposed framework of training an VLMM with AI feedback as Video Large Multimodal model with RLAIF or VLM-RLAIF for short.
 Our empirical studies show that our aligned VLM-RLAIF exhibits superior performance compared to state-of-the-art VLMMs across a wide array of video benchmarks, as illustrated in Fig.
@@
 Multimodal large model.
-In recent research efforts, there has been a significant focus on advancing Language Models (LLMs) to go beyond natural language understanding, extending into the realm of multi-modal comprehension.
-The goal is to develop LLMs capable of interpret additional modalities, e.g., image Liu et al.
+Recently, there have been significant advances for LLMs to go beyond natural language understanding, extending into the realm of multimodal comprehension.
+The goal is to develop LLMs capable of understanding various modalities, e.g., image Liu et al.
 (2023a), video Li et al.
-(2023a); Lin et al.
+(2023b); Lin et al.
 (2023), 3D point-cloud Guo et al.
-(2023), etc.
-To provide additional modality into LLMs, most of work utilize pre-trained encoder, such as CLIP Radford et al.
-(2021), Q-former Li et al.
+(2023) and etc.
+To make the LLMs multimodal, most of the work utilize a pretrained encoder, such as CLIP Radford et al.
+(2021a), Q-former Li et al.
 (2022) or ImageBind Girdhar et al.
-(2023), to extract each modality feature from data.
-These features then projected to the token embedding space of the language model.
-Following this, the models are subject to supervised fine-tuning (SFT) using synthetically generated, modality-specific instruction-following datasets.
-These approaches, exemplified in works such as LLaVA Liu et al.
+(2023), to extract each modality’s representations from data.
+These representations are then projected into the token embedding space of the language model.
+Then, the models undergo supervised fine-tuning (SFT) with synthetically generated, modality-specific instruction-following datasets.
+These approaches, adopted in LLaVA Liu et al.
 (2023a), Video-LLaVA Lin et al.
 (2023) or Point-LLM Guo et al.
-(2023), facilitates the development of models proficient in engaging in conversations grounded in additional modality.
-Figure 3: Detailed illustration of the three stages for context-aware reward modeling.
-Dotted box in each stage represents distinct model input: the first is for generating responses using the SFT model, the second for the judge model to evaluate and choose the superior response between options A and B, and the third for training the RM.
-Each input includes a task-specific prompt, denoted by the yellow box, tailored to guide the model’s operation within its respective function (more details are in Appendix Sec.
-C).
-The operational flow of each stage is as follows: (1) The SFT model produces two candidate responses from the provided video and question.
-(2) With the video, question and responses at hand, the SFT model utilize context information and guiding prompt to evaluate the responses.
-(3) The RM is trained using the preference pairs generated in the previous step as indicated in orange box.
-Figure 4: An example of context-aware preference selection.
-We demonstrate our model’s ability to generate preference feedback, i.e., preferred choice and explanation marked in orange box, on given instruction input, two responses and context.
-Red color denotes an incorrect response, while blue color signifies a correctly grounded response with respect to the visual input.
-The rationale behind selecting ‘Output B’ as the preferred choice is indicated in green.
+(2023), facilitate the development of proficient conversations grounded in additional modality.
 Reinforcement learning from feedback.
@@
 (2023a).
-By collecting the preferences from human evaluators, it usually trains the reward model that gives a high reward to the preferred output of the model.
-However, significant challenge in this process is the annotation cost associated with selecting preference.
-To mitigate this issue, the concept of Reinforcement Learning with AI-inferred Feedback (RLAIF) was introduced Bai et al.
+By collecting preferences from human evaluators, it usually trains the reward model that gives a high reward to the preferred output of the model.
+However, a significant challenge in this process is the annotation cost associated with selecting the preference.
+To mitigate this issue, Reinforcement Learning from AI Feedback (RLAIF) was proposed Bai et al.
 (2022); Lee et al.
@@
 (2023b).
-RLAIF capitalizes on the inherent ability of Large Language Models (LLMs) to rate or evaluate information, allowing the LLM itself to assign preferences.
+RLAIF capitalizes on the inherent ability of Large Language Models (LLMs) to evaluate the generated responses from the SFT model, allowing the LLM itself to assign preferences.
 3 VLM-RLAIF Framework
-3.1 Overview
-Contrasting to the limited scalability of human feedback in RLHF, we leverage AI’s feedback for aligning multimodality between video and text, reducing the reliance on exhaustive human-annotated preferences Ouyang et al.
+To overcome the limited scalability of human feedback in RLHF, we use AI’s feedback to align multimodality between video and text, reducing the reliance on exhaustive human-annotated preferences Ouyang et al.
 (2022); Sun et al.
 (2023a).
+In particular, we improve the feedback process by using detailed video descriptions, thereby achieving better contextual clarity in video content.
 The training procedure of VLM-RLAIF can be summarized into three stages as follows:
 Supervised fine-tuning (SFT).
-We fine-tune an LLM, i.e., Vicuna, using supervised learning on synthetically generated video-text instruction-tune data Muhammad Maaz and Khan (2023).
-This involves integrating a vision encoder with two linear layers and an additional learnable parameters using LoRA Hu et al.
+We first fine-tune an LLM, e.g., Vicuna, using supervised learning on synthetically generated video-text instruction-tune data Muhammad Maaz and Khan (2023).
+This involves the integration of a vision encoder with two linear layers and additional learnable parameters using LoRA Hu et al.
 (2022), into the training process.
-Such fine-tuning enhances the model’s capability to follow instructions.
-We refer to this model as the Video Large Multimodal model with SFT or VLM-SFT for short.
+This fine-tuning allows the model to better follow the instructions Muhammad Maaz and Khan (2023); Su et al.
+(2023).
+Additionally, we improve the SFT process by expanding the instruction-tune data and introducing simple curriculum learning (Sec.
+3.2).
+We refer to this fine-tuned model as the Video Large Multimodal model with SFT or VLM-SFT for short.
 Reward modeling with AI feedback.
@@
 (2023).
-To accomplish this, we employ the VLM-SFT as a judge for assessing preferences.
-Once preferences are annotated, we proceed to train a reward model (RM) based on these annotated labels using a cross-entropy loss, following the Bradley-Terry model for estimating score functions from pairwise preferences Ouyang et al.
+To obtain human-like preference, we employ the VLM-SFT as a judge to assess preferences.
+Once preferences are judged, we train a reward model (RM) based on preferences using a cross-entropy loss, following the Bradley-Terry model for estimating score functions from pairwise preferences Ouyang et al.
 (2022); Sun et al.
 (2023a).
-We illustrate the training procedure for collecting preferences and training the reward model in the Fig.
-3.
-The RM give higher score reward to the better response and lower score reward to the less appropriate one in a pair of responses, thus providing guidance to a policy model during the reinforcement learning (RL) process.
-Reinforcement learning.
-During the RL phase, we fine-tune a supervised policy model, initialized from the VLM-SFT, aiming to optimize the scalar reward output from the trained RM.
-We use the Proximal Policy Optimization (PPO) algorithm Schulman et al.
-(2017) for this purpose, following prior research Ouyang et al.
+We describe the training procedure for collecting preferences and training the reward model in Sec.
+3.1.
+The RM give higher score reward to the better response and lower score reward to the less appropriate one in a pair of responses (see examples in Appendix Fig.
+12), thus guiding the policy model using reinforcement learning (RL).
+Reinforcement learning from AI feedback.
+We finally fine-tune a supervised policy model, initialized from the VLM-SFT, aiming to optimize the scalar reward output of the trained RM by reinforcement learning.
+Specifically, we use the Proximal Policy Optimization (PPO) algorithm Schulman et al.
+(2017), following Ouyang et al.
 (2022); Sun et al.
 (2023b, a).
-This process finely tunes the model to respond optimally to the rewards, thereby enhancing its decision-making abilities in accordance with the learned preferences.
 3.1 Context-Aware Reward Modeling
-For effective utilization of VLM-SFT in preference selection, a nuanced understanding of video content is essential, particularly in scenarios involving complex scenes with elements subject to varied interpretations.
-Acknowledging this intricacy, it becomes apparent that exclusive dependence on VLM-SFT for direct video analysis may fall short in achieving thorough comprehension.
+For VLM-SFT to select preference grounded on the video, we argue that a detailed understanding of video content is necessary for more accurate and contextually relevant decisions by the VLM-SFT.
+However, the current video encoder presents challenges in accurately encoding the temporal details of videos as they are based on the image encoder Radford et al.
+(2021b).
+Figure 3: The three stages of the proposed context-aware reward modeling.
+The work flow of each stage is as follows: (1) The SFT model produces two candidate responses from the provided video and question.
+(2) With the video, question and responses at hand, the SFT model utilize context information and guiding prompt to evaluate the responses.
+(3) The RM is trained using the preference pairs generated in the previous step as indicated in orange box.
+Dotted box in each stage denotes a model’s input: the first is for generating responses using the SFT model, the second is for the judge model to evaluate and choose the superior response between options A and B, and the third is for training the RM.
+Each input includes a task-specific prompt, denoted by the yellow box, tailored to guide the model’s operation within its respective function (more in Appx.
+Sec.
+B).
+Figure 4: An example of context-aware preference selection.
+We demonstrate our model’s ability to generate preference feedback, i.e., preferred choice and explanation marked in orange box, on given instruction input, prompt, two responses and context.
+Red color denotes an incorrect response, while blue color signifies a correctly grounded response with respect to the visual input.
+The rationale behind selecting ‘Output B’ as the preferred choice is indicated in green.
 Context-aware preference selection.
-In response, we enhance the VLM-SFT by integrating detailed video descriptions, referred to as context, into the preference selection workflow, thereby imparting additional contextual clarity, as illustrated in Figures 3-(2) and 4.
-This process entails breaking down the video into smaller clips, followed by generating descriptive caption for each segment to obtain fine-grained information.
-Subsequently, these individual captions are concatenated into an exhaustive narrative of the video.
-This detailed narrative is then provided to the judge model, i.e., VLM-SFT, to aid in better preference selection.
-This process not only improves the VLM-SFT’s ability to obtain a comprehensive view of the video content, but also enables it to more identify the most suitable response.
-Integrating all inputs, i.e., marked in dotted box in Fig.
-3-(2), with a specific prompt, i.e., rules for generating preferences as illustrated in Appendix Fig.
-9, allows us to collect context-aware synthetic preferences, thus enriching the model’s decision-making process.
-Training reward model.
-The reward model (RM) is designed to assign higher scores to responses considered better and lower scores to those considered lower in quality.
-Starting from VLM-SFT with the final linear layer removed, we train a model to take in an input prompt and response, i.e., marked in dotted box of Fig.
+We propose to explicitly integrate detailed video descriptions, referred to as context, into the preference selection workflow, thereby imparting additional contextual clarity to the VLMM, as illustrated in Figures 3-(2) and 4.
+Specifically, we start by segmenting the video into small clips, each containing up to 20 frames, and then employ the VLM-SFT to generate a detailed video description for each segment with an input prompt, ‘Describe this video in detail’.
+Subsequently, these individual captions are concatenated, which we call a narrative of the video.
+The narrative is then provided to a judge model, i.e., VLM-SFT, for better preference selection.
+The context not only improves the VLM-SFT’s ability to obtain a comprehensive view of the video content, but also enables it to identify the most suitable response for the video (see Sec.
+4.3 for empirical results).
+Integrating the context with instruction inputs with a specific prompt (rules for generating preferences as illustrated in Appendix Fig.
+9), marked in dotted boxes in Fig.
+3-(2), allows us to collect context-aware preferences.
+Training the reward model.
+We design the reward model (RM) to assign higher scores to responses considered better and lower scores to those considered worse in quality.
+Starting from VLM-SFT with the final linear layer removed, we train a model to take an input prompt and response, i.e., marked in dotted boxes of Fig.
 3-(3), and output a scalar reward value.
 Using the preference dataset produced from the VLM-SFT, we train the RM with a cross-entropy loss.
-Specifically, we use 13B VLM-SFT to train the reward model, as it gives slightly better performance than 7B VLM-SFT (see the quantitative comparison discussed in Appendix Sec.
-A).
-Remarkably, after training RLAIF using the RM, the 7B VLM-RLAIF significantly surpass the 13B VLM-SFT, thus validating the effectiveness of our proposed framework in aligning video and text modalities (refer to Sec.
+Specifically, we use 13B VLM-SFT to train the reward model, as it gives slightly better performance than 7B VLM-SFT (see the quantitative comparison discussed in Sec.
+4.3).
+Note that, after training RLAIF using the RM, the 7B VLM-RLAIF significantly surpass the 13B VLM-SFT, thus validating the effectiveness of our proposed framework in aligning video and text modalities (refer to Sec.
 4.2 for more details).
-3.3 Two-stage Curriculum Supervised Fine-Tuning
-To obtain video conversational capability, we train the LLM with a open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023).
-Further, to enhance this, we not only expand our training with additional video-text instruction-tune datasets but also introduce a simple curriculum learning strategy.
-This approach is designed to improve the model’s performance progressively, ensuring a more effective and nuanced understanding of video-text interactions.
+3.2 Two-stage Curriculum SFT
+During the SFT process, we initially train the LLM with a open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023).
+To improve the VLMM, we not only augment our training with additional video-text instruction-tune datasets but also propose a novel curriculum learning strategy.
 Augmenting video instruction-tune dataset.
-To improve the video understanding ability, we first augment the open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023) with existing human-annotated video question answering datasets Xiao et al.
+To improve the video understanding ability, we first augment the video-text instruction-tune dataset Muhammad Maaz and Khan (2023) with existing human-annotated video question answering datasets Xiao et al.
 (2021); Li et al.
 (2020).
-Moreover, recognizing the importance of aligning LLMs with the video modality, we focus on obtaining instruction-tune dataset that encompass both visual and temporal intricacies.
-To achieve this, we leverage Video Localized Narratives Voigtlaender et al.
+In particular, we focus on obtaining instruction-tune dataset that encompass both visual and temporal intricacies for video comprehension.
+To obtaining such dataset, we use Video Localized Narratives Voigtlaender et al.
 (2023), a dataset that provides detailed descriptions of the appearance and complex actions of all objects in the video.
-We transform the narrative dataset into object-centric instruction-tune dataset format using ChatGPT similar to prior work Muhammad Maaz and Khan (2023).
-More details about the generated instruction-tune dataset can be found in the Appendix Sec.
-B.
-We train the SFT model using both the open-sourced and our proposed datasets, resulting in improved performance, as discussed in the Sec.
-4.3.
-Curriculum supervised fine-tuning
-Recent VLMM research has advanced instruction-following capabilities through curating multimodal instruction-tune data Muhammad Maaz and Khan (2023); Luo et al.
-(2023), joint learning across image and video datasets Lin et al.
-(2023); Li et al.
-(2023c), and enhanced bridging networks like adapter Liu et al.
-(2023b).
-Unlike these approaches, we emphasize a structured learning method for VLMMs, inspired by the human learning process.
-We advocate for the importance of the training sequence, drawing from curriculum learning principles Chang et al.
-(2021), and propose a simple two-stage curriculum learning strategy for SFT, aiming to optimize the instruction-following proficiency of VLMMs.
-To accomplish this, we divide the SFT dataset based on complexity into two segments: ‘easy’ and ‘hard’.
+We transform the narrative dataset into an object-centric instruction-tune dataset format using ChatGPT similar to prior work Muhammad Maaz and Khan (2023).
+More details about the generated instruction-tune dataset are in Appendix Sec.
+A.
+Curriculum-based supervised fine-tuning.
+Inspired by the human learning process, we argue the importance of the training sequence, the curriculum learning principles Chang et al.
+(2021) for learning VLMMs.
+We propose a simple two-stage curriculum learning strategy for SFT, aiming to optimize the instruction-following proficiency of VLMMs.
+For the curriculum based structured learning, we divide the SFT dataset based on complexity into two segments: ‘easy’ and ‘hard’.
 The model is first trained on ‘easy’ tasks to learn basic concepts, then progressed to ‘hard’ tasks for advanced challenges.
-Given that longer answers often require deeper comprehension of the context and enhanced proficiency in creating syntactically complex sentences Xu et al.
+For the difficulty or easiness of the data, given that longer answers often require deeper comprehension of the context and enhanced proficiency in creating syntactically complex sentences Xu et al.
 (2020); Agrawal and Singh (2023); Ranaldi et al.
-(2023), we use answer length as our proxy for sample difficulty, meaning that the longer the answer sentence, the more difficult the task is.
-This proxy also allows for a clear basis for dividing data, offering a concrete measure of difficulty that aligns the model’s learning trajectory with the incremental nature of human learning.
+(2023), we use answer length as our criterion for sample difficulty, i.e., the longer the answer sentence, the more difficult the task is.
 Methods LLM Size Video-based Generative Performance
-Correctness Detail Context Temporal Consistency
-VideoChat Li et al. (2023a) 7B 2.23 2.50 2.53 1.94 2.24
+Correctness $\uparrow$ Detail $\uparrow$ Context $\uparrow$ Temporal $\uparrow$ Consistency $\uparrow$
+VideoChat Li et al. (2023b) 7B 2.23 2.50 2.53 1.94 2.24
 LLaMA-Adapter Zhang et al. (2023b) 7B 2.03 2.32 2.30 1.98 2.15
@@
 VideoChat2 Li et al. (2024) 7B 3.02 2.88 3.51 2.66 2.81
-LLaMA-VID Li et al. (2023c) 7B 2.96 3.00 3.53 2.46 2.51
-LLaMA-VID Li et al. (2023c) 13B 3.07 3.05 3.60 2.58 2.63
+LLaMA-VID Li et al. (2023d) 7B 2.96 3.00 3.53 2.46 2.51
+LLaMA-VID Li et al. (2023d) 13B 3.07 3.05 3.60 2.58 2.63
 GPT-4V OpenAI et al. (2023) - 3.85 3.45 3.84 3.63 2.8
@@
 $\Delta$ (RLAIF - SFT) - +0.84 +0.43 +0.63 +0.95 +0.83
-Table 1: Qualitative comparison between different VLMMs on video-based generative performance benchmark.
-We assess various VLMMs on video-based generative benchmark which measures various generative aspects.
-Our approach, VLM-RLAIF, shows a performance improvement over previous approaches, with the exception of GPT-4V, and demonstrates noticeable enhancements across five criteria when compared to the VLM-SFT.
+Table 1: Quantitative comparison between different VLMMs on video-based generative performance benchmark.
+Our approach, VLM-RLAIF, shows a performance improvement over previous approaches, with the exception of GPT-4V which requires much more computational resource than ours, and demonstrates noticeable enhancements across five criteria when compared to the VLM-SFT.
 Here, $\Delta$ (RLAIF - SFT) indicates the improvement of RLAIF model over SFT model.
@@
 4.1 Experimental Setup
-Model settings.
-We initiate training of the VLM-SFT, based on the pre-trained LLaVA image-text large multimodal models Liu et al.
-(2023a), leveraging a diverse video-text instruction-tune dataset Muhammad Maaz and Khan (2023); Xiao et al.
-(2021); Li et al.
-(2020); Voigtlaender et al.
-(2023).
-This training incorporates a curriculum approach as outlined in Sec.
-3.3.
-Upon establishing the VLM-SFT, we proceed to train the reward model (RM) using the VLM-SFT for both its foundation and the generation of preference data.
-Subsequently, we implement reinforcement learning with the trained RM, leading to the development of the VLM-RLAIF.
-In the reinforcement learning phase, the policy model is initialized from the VLM-SFT, while the value model employs the RM.
+Model details.
+We initiate training of the VLM-SFT, building on a pre-trained image-text model Liu et al.
+(2023a), with various video-text instruction-tune datasets.
+In particular, we employ a video projection layer consisting of two linear layers with ReLU activation function between them.
+Upon establishing the VLM-SFT, we train the RM using the VLM-SFT for both its foundation and the generation of preference data.
+Subsequently, we train the RM using reinforcement learning (VLM-RLAIF).
+The policy model is initialized from the VLM-SFT, while the value model is initialized from the RM.
 Dataset details.
-For the SFT dataset, we utilize the open-sourced video-text instruction-tune dataset (80k) Muhammad Maaz and Khan (2023); Li et al.
-(2023a) and video question answering datasets (67k) Xiao et al.
+For the SFT dataset, we utilize the open-source video-text instruction-tune dataset (80k) Muhammad Maaz and Khan (2023); Li et al.
+(2023b) and video question answering datasets (67k) Xiao et al.
 (2021); Li et al.
 (2020).
-Additionally, we generate object-centric narrative video-text instruction-tune dataset (180k) for training the VLM-SFT, as mentioned in Sec.
-3.3.
-For two-stage curriculum learning, we divide the instruction-tune data for SFT into two groups based on the difficulty; easy (214k) and hard (113k) data.
-To train the reward model (RM), we first generate responses from existing instruction-tune data Muhammad Maaz and Khan (2023) and generate preferences using them (40k).
-Then, we once again utilize the existing instruction-tune dataset (100k) Muhammad Maaz and Khan (2023) for reinforcement learning with the trained RM for initialization.
-Evaluation details.
-We evaluate our VLM-RLAIF on various video benchmarks; zero-shot video questions and video-based generative benchmarks following the prior work Muhammad Maaz and Khan (2023); Lin et al.
-(2023); Li et al.
-(2023c) and text-to-video-retrieval and action recognition as proposed in prior work Li et al.
-(2023b).
+More importantly, we generate object-centric narrative video-text instruction-tune dataset (180k) for training the VLM-SFT (Sec.
+3.2).
+For the two-stage curriculum learning, we divide the instruction-tune data for SFT into two groups based on the difficulty; easy (214k) and hard (113k) data.
+To train the RM, we first generate responses from existing instruction-tune data Muhammad Maaz and Khan (2023) and generate preferences using them (40k).
+Then, we again use the existing instruction-tune dataset (100k) Muhammad Maaz and Khan (2023) for RL with the trained RM.
 Training details.
@@
 FrozenBiLM Yang et al. (2022) 1B 32.2 - 16.8 - 24.7 -
-VideoChat Li et al. (2023a) 7B 56.3 2.8 45.0 2.5 26.5 2.2
+VideoChat Li et al. (2023b) 7B 56.3 2.8 45.0 2.5 26.5 2.2
 LLaMA-Adapter Zhang et al. (2023b) 7B 54.9 3.1 43.8 2.7 34.2 2.7
@@
 VideoChat2 Li et al. (2024) 7B 70.0 3.9 54.1 3.3 49.1 3.3
-LLaMA-VID Li et al. (2023c) 7B 69.7 3.7 57.7 3.2 47.4 3.3
-LLaMA-VID Li et al. (2023c) 13B 70.0 3.7 58.9 3.2 47.5 3.3
+LLaMA-VID Li et al. (2023d) 7B 69.7 3.7 57.7 3.2 47.4 3.3
+LLaMA-VID Li et al. (2023d) 13B 70.0 3.7 58.9 3.2 47.5 3.3
 VLM-SFT 7B 67.2 3.6 52.4 3.0 44.1 3.2
@@
 $\Delta$ (RLAIF - SFT) - +9.2% +0.4 +10.6% +0.4 +13.2% +0.3
-Table 2: Qualitative comparison between different VLMMs on zero-shot video question answering benchmark. Our VLM-RLAIF outperforms previous work across three video-question answering benchmarks.
+Table 2: Quantitative comparison between different VLMMs on zero-shot video question answering benchmark. VLM-RLAIF outperforms previous work across three video-question answering benchmarks.
 Methods LLM Size T2V Retrieval Action Recognition
@@
 Video-LLaVA Lin et al. (2023) 7B 29.34 55.35 18.70 38.60 52.33 80.86 36.64 64.03
-LLaMA-VID Li et al. (2023c) 7B 27.28 53.40 17.00 35.10 56.58 82.79 38.85 65.27
+LLaMA-VID Li et al. (2023d) 7B 27.28 53.40 17.00 35.10 56.58 82.79 38.85 65.27
 VLM-SFT 7B 26.65 54.27 13.10 30.50 53.03 80.34 38.58 62.37
@@
 $\Delta$ (RLAIF - SFT) - +9.38 +9.13 +7.90 +10.2 +9.80 +5.52 +8.11 +6.00
-Table 3: Qualitative comparison between different VLMMs on zero-shot text-to-video (T2V) retrieval and action recognition.
-Following prior work Li et al.
-(2023b), we assess our proposed VLM-RLAIF on zero-shot T2V and action recognition for comprehensive evaluation of the VLMMs.
-$\dagger$} denotes reproduced results using the author’s implementation.
-Figure 5: Effect of various preference data size on video-based generative benchmark. VLM-RLAIF performance increase with the number of collected preference data.
+Table 3: Quantitative comparison between different VLMMs on zero-shot text-to-video (T2V) retrieval and action recognition.
+Following Li et al.
+(2023c), we evaluate our proposed VLM-RLAIF on zero-shot T2V retrieval and action recognition.
+$\dagger$}: reproduced by the authors’ implementation.
+Figure 5: Effect of preference data size on video-based generative benchmark.
+VLM-RLAIF’s performance improves across five metrics as the amount of collected preference data increases.
+The metrics evaluate whether the model generates correct descriptions of the video, includes detailed explanations, remains contextual, demonstrates temporal understanding, and provides consistent responses to variations of the same question.
 4.2 Quantitative Analysis
-We quantitatively evaluate our proposed VLM-RLAIF on various video benchmarks following prior work: video-based generative benchmark, zero-shot video question answering Muhammad Maaz and Khan (2023); Lin et al.
+We evaluate our proposed VLM-RLAIF on various video benchmarks including video-based generative benchmark, zero-shot video question answering Muhammad Maaz and Khan (2023); Lin et al.
 (2023); Li et al.
-(2023c), text-to-video retrieval and action recognition Li et al.
-(2023b).
+(2023d), text-to-video retrieval, and action recognition Li et al.
+(2023c).
 Video-based generative performance.
-We evaluate VLMMs on the video-based generative performance benchmark Muhammad Maaz and Khan (2023) that measures five criteria: correctness, detail, context, temporal understanding, and consistency of generated text.
+We evaluate VLMMs on the video-based generative performance benchmark Muhammad Maaz and Khan (2023) that measures five criteria of generated text.
 In specific, these assess the relevance of the model’s output to the video content, its capacity to capture essential details and contextual information, its understanding of temporal sequences, and the consistency in responding to varied yet related queries.
-As indicated in Tab.
+As shown in Tab.
 1, the VLM-RLAIF performs on par with GPT-4V OpenAI et al.
-(2023) and outperforms previous approaches and the VLM-SFT.
+(2023), which requires much more computational resources than ours (i.e., not a fair comparison), and outperforms previous approaches and the VLM-SFT.
 Zero-shot video question answering.
-To assess the general capabilities of various video-text Large Language Models (LLMs), we conduct a quantitative evaluation of their video question answering (VideoQA) abilities using three datasets: MSVD-QA, MSRVTT-QA Xu et al.
-(2017), and ActivityNet-QA Yu et al.
-(2019).
-We adopted the evaluation pipeline from prior work Muhammad Maaz and Khan (2023), which utilizes ChatGPT-assisted evaluation, to gauge the zero-shot VideoQA performance.
-The results, as shown in Table 2, indicate that the VLM-RLAIF significantly outperforms previous models.
+To evaluate the reasoning ability of VLLMs, we conduct a quantitative evaluation of video question answering (VideoQA) abilities on three datasets Xu et al.
+(2017); Yu et al.
+(2019), following Muhammad Maaz and Khan (2023).
+The results, as shown in Table 2, indicate that the VLM-RLAIF significantly outperforms previous approaches, including VLM-SFT.
 Notably, VLM-RLAIF exceeds VLM-SFT by 9.2%, 10.6%, and 13.2% in accuracy and by 0.4, 0.4, and 0.3 in score across all datasets.
+We believe that the better visually-aligned response generated from the VLM-RLAIF improves the performance (see quantitative analysis in Figures 6 and 11).
 Zero-shot text-to-video retrieval.
-For this evaluation, we follow the procedure proposed in prior work Li et al.
-(2023b), which compare CLIP score between generated description and ground-truth caption.
+For this task, we follow the procedure proposed in Li et al.
+(2023c), which compares CLIP score between generated description and ground-truth caption.
 Table 3 illustrates the summarized performance comparison to various VLMMs.
-In two datasets, i.e., MSVD and MSRVTT, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
+In the two datasets, i.e., MSVD and MSRVTT, the proposed VLM-RLAIF clearly outperforms other methods including our VLM-SFT by the help of better alignment by the proposed components.
 Zero-shot action recognition.
-Following the VLMMs evaluation procedure proposed in prior work Li et al.
-(2023b), we conduct zero-shot action recognition task using two benchmark datasets, e.g., UCF101 and HMDB51.
-Table 3 illustrates the summarized performance comparison to various VLMMs.
-In two datasets, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
+Following the VLMMs evaluation procedure proposed in Li et al.
+(2023c), we conduct zero-shot action recognition task using two benchmark datasets, e.g., UCF101 and HMDB51.
+We summarize results of various VLMMs in Tab.
+3.
+In two datasets, the proposed VLM-RLAIF again clearly outperforms other methods including our VLM-SFT.
 SFT datasets Curr. learning Video-based Generative Performance
-[A] [B] [C] Corr. Det. Cont. Temp. Cons.
+[A] [B] [C] Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
 2.32 2.53 3.03 2.16 2.23
@@
 ‘[A]’ indicate the multimodal instruction-tune dataset proposed in Muhammad Maaz and Khan (2023); Li et al.
-(2023a).
+(2023b).
 ‘[B]’ represents the use of a human-labeled video question answering dataset Xiao et al.
@@
 (2020), while ‘[C]’ refers to the use of an object-centric video narrative instruction-tune dataset (Appendix Sec.
-B).
+A).
 ‘Curr.
-learning’ indicates the application of curriculum learning as detailed in Sec.
-3.3.
+learning’ indicates the curriculum learning (Sec.
+3.2).
 RLAIF Context #\#Clips Video-based Generative Performance
-Info. Corr. Det. Cont. Temp. Cons.
+Info. Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
 - 2.79 2.82 3.37 2.28 2.49
@@
 3 3.63 3.25 4.00 3.23 3.32
-Table 5: Effect of context information on video-based generative performance benchmark Muhammad Maaz and Khan (2023).
+Table 5: Effect of context information on video-based generative performance benchmark.
 We investigate the efficacy of using context information for reward modeling (Sec.
-3.2).
+3.1).
 ‘Context Info.’ indicates the use of context in preference selection.
-‘#\# Clips’ refers to the the number of times we divide the video into segments to generate the context information.
+‘#\# Clips’ denotes the number of segments into which we divide the video to generate the context information.
+Methods LLM Size Video-based Generative Perf.
+Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
+VLM-SFT 7B 2.32 2.53 3.03 2.16 2.23
+VLM-SFT 13B 2.64 2.73 3.28 2.38 2.44
+Table 6: Quantitative comparison between different sizes of the VLM-SFT.
+We assess the performance of VLM-SFT with varying LMM sizes, specifically 7B and 13B, on video-based generative benchmarks.
+We conduct this evaluation without the integration of augmented instruction-tune data and the implementation of two-stage curriculum learning.
+Methods LLM Size Video Question Answering Video-based Generative Perf.
+Policy Reward MSVD MSRVTT ActivityNet Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
+Model Model Acc. Score Acc. Score Acc. Score
+LLaMA-VID (7B) - - 69.7 3.7 57.7 3.2 47.4 3.3 2.96 3.00 3.53 2.46 2.51
+VLM-SFT 7B - 67.2 3.6 52.4 3.0 44.1 3.2 2.79 2.82 3.37 2.28 2.49
+VLM-RLAIF 7B 7B 75.1 3.9 61.0 3.3 56.1 3.4 3.47 3.14 3.87 3.05 3.30
+VLM-RLAIF 7B 13B 76.4 4.0 63.0 3.4 57.3 3.5 3.63 3.25 4.00 3.23 3.32
+Table 7: Quantitative comparison between different size of policy model and reward model for the VLM-RLAIF.
+We evaluate the VLM-RLAIF with different model size for policy model and reward model on zero-shot video question answering and video-based generative benchmark.
 4.3 Detailed Analysis
-For a detailed analysis, we leverage the video-based generative benchmark Muhammad Maaz and Khan (2023) specifically, as it is well suited to assess the wide-ranging capabilities of VLMMs.
+For a detailed analysis, we use the video-based generative benchmark Muhammad Maaz and Khan (2023) specifically, as it is well suited to evaluate the wide-ranging capabilities of VLMM, i.e., focusing on response relevance, detail and context capture, temporal understanding, and consistency across queries.
 In-depth analysis of SFT training.
 We first empirically support the effectiveness of augmenting the SFT dataset with additional instruction-following dataset (Sec.
-3.3).
+3.2).
 The first and second rows of Tab.
-4 illustrate the benefits of incorporating this additional dataset in improving video-based generative performance.
-On top of that, the application of curriculum learning significantly boosts performance, demonstrating the efficacy of employing a curriculum learning strategy during the SFT process, as illustrated in the third row of Tab.
-4.
-Effect of various preference data size.
-Our framework’s strength lies in generating synthetic preference feedback in large quantities.
-To evaluate this capability, we escalated the amount of preference data from 10k – the same quantity utilized in the multimodal image-text RLHF framework Sun et al.
+4 illustrate the benefits of incorporating this additional dataset in improving performance.
+On top of that, the application of curriculum learning significantly improves performance, implying the efficacy of curriculum learning for the SFT process (the third row of Tab.
+4).
+Effect of preference data size.
+Our method’s strength lies in generating synthetic preference feedback in large quantities.
+To study the benefit of large-sized data, we sweep the size of preference data from 10k – the same quantity utilized in the multimodal image-text RLHF framework Sun et al.
 (2023a) – to 40k.
-Figure 5 shows a steady improvement in performance with the increase in data size.
+As expected, we observe monotonic increases in performance by the increase in data size, as shown in Figure 5.
+Methods LLM Size T2V Retieval Action Recognition
+Policy Reward MSVD MSRVTT UCF101 HMDB51
+Model Model R@1 R@5 R@1 R@5 Top-1 Top-5 Top-1 Top-5
+LLaMA-VID (7B) - - 27.28 53.40 17.00 35.10 56.58 82.79 38.85 65.27
+VLM-SFT 7B - 26.65 54.27 13.10 30.50 53.03 80.34 38.58 62.37
+VLM-RLAIF 7B 7B 33.73 61.95 20.80 42.90 61.09 85.15 43.86 65.88
+VLM-RLAIF 7B 13B 36.03 63.40 21.00 40.70 62.83 85.86 44.75 68.37
+Table 8: Quantitative comparison between different size of policy model and reward model for the VLM-RLAIF.
+We evaluate the VLM-RLAIF with different model size for policy model and reward model on zero-shot text-to-video retrieval and action recognition tasks.
+Figure 6: Samples of VLM-RLAIF and VLM-SFT in video understanding.
+Two samples demonstrate better capability of the VLM-RLAIF.
+Red color indicates visually un-grounded expressions considering video content and blue color indicates well grounded expressions to the video.
 Context-aware preference labeling.
-Table 5 presents a study on the impact of context information on the performance of the video-based generative benchmark discussed in Sec.
-3.2.
-This reveals that the use of context information during preference selection improves model performance, as demonstrated in the second and third rows in Table 5.
-Moreover, segmenting the video into multiple clips, specifically three, and using their aggregated caption as context information further enhances performance, as indicated in the third and fourth rows.
-We conjecture that segmenting video into clips and providing detailed captions for each clip enables the model to capture more detailed nuances, thereby enriching the contextual clarity for reward modeling and boosting the VLM-RLAIF’s performance.
-Figure 6: Samples of VLM-RLAIF and VLM-SFT in video understanding.
-Two samples demonstrate superior capability of the VLM-RLAIF.
-Red color indicates visually un-grounded expressions considering video content and blue color well grounded to the visual content.
+Table 5 presents a study on the impact of context information.
+This reveals that the use of context information during preference selection improves performance, as demonstrated in the second and third rows in Tab.
+5.
+Moreover, segmenting the video into multiple clips, specifically three, and using their aggregated captions as context further improves performance (see the third and fourth rows).
+We conjecture that segmenting video into clips with detailed captions enables the model to capture detailed nuances, thereby enriching the contextual clarity for reward modeling.
+Comparison between different SFT model sizes.
+Table 6 shows a clear performance margin between VLM-SFT with different LLM sizes, specifically 7B and 13B, in video-based generative tasks.
+Larger LLM sizes lead to improved performance, suggesting that increased model capacity enhances the ability to capture and generate complex video content.
+Thus, we adopt the 13B model for the reward model (Sec.
+3.1), with supporting evidence in Tables 7 and 8.
+Various LLM size for reward model
+Tables 7 and 8 show the performance of VLM-RLAIF with different RM sizes, initialized from VLM-RLAIF-7B and VLM-RLAIF-13B.
+In both cases, the policy model is initialized with VLM-SFT-7B.
+Our RLAIF method outperforms VLM-SFT significantly across all evaluation methods.
+Specifically, RLAIF with 7B RM achieves a 5-12% improvement in zero-shot video question answering.
+Scaling up the RM from 7B to 13B further improves performance, except for the text-to-video retrieval task R@5 metric on MSR-VTT.
 4.4 Qualitative Analysis
-VLM-SFTvs. VLM-RLAIF.
-Figure 6 qualitatively compares the performance of VLM-SFT and VLM-RLAIF across two video samples, highlighting their multimodal understanding capabilities.
-VLM-RLAIF consistently yields more accurate answers than VLM-SFT, as seen in the detailed recognition of an athlete’s attire in the first example, marked in red and blue.
-The second example further affirms VLM-RLAIF’s superiority in generating precisely grounded responses to the visual input, where VLM-SFT falls short.
+We now qualitatively compares the performance of VLM-SFT and VLM-RLAIF, highlighting their multimodal understanding capabilities in Figure 6.
+VLM-RLAIF consistently yields more accurate answers than VLM-SFT, as shown in the detailed recognition of the attire of an athlete in the first example, marked in red and blue.
+The second example further affirms VLM-RLAIF’s benefit in generating better grounded responses to the visual input, where VLM-SFT falls short.
 More examples are in the Appendix Fig.
-11.
+11 for the space sake.
 5 Conclusion
-We introduce a novel alignment strategy for VLMMs, termed VLM-RLAIF, utilizing reinforcement learning from AI feedback.
-To enhance multimodal alignment, we also present a context-aware reward modeling technique, enabling AI to generate feedback for self-improvement.
-Additionally, we expand the instruction-tune dataset for SFT and adopt a straightforward curriculum training approach.
-The VLM-RLAIF significantly surpasses previous models across multiple multimodal video-text understanding benchmarks, including video-based generative tasks, zero-shot VideoQA, text-to-video retrieval, and action recognition.
-Appendix A Quantitative Comparison between Different SFT Model Sizes
-We assess the performance of the VLM-SFT using varying LLM sizes, specifically 7B and 13B, across video question answering (VideoQA) and video-based generative benchmarks, detailed in Table 6.
-The results indicate similar outcomes in the zero-shot VideoQA tasks across both model sizes.
-However, in video-based generative tasks, a clear trend emerges: larger LLM sizes lead to improved performance.
-This suggests that the increased model capacity inherent in larger LLMs enhances the model’s ability to capture and generate more complex and nuanced video content representations.
-Thus, we adopt the 13B model for the reward model (Sec.
-3.2), with empirical evidence provided in Tables 7 and 8.
-Appendix B Details About Object-Centric Instruction-Tune Data
+We propose a novel alignment strategy for VLMMs, termed VLM-RLAIF, that uses reinforcement learning from AI feedback.
+To improve multimodal alignment, we propose a context-aware reward modeling, enabling AI to generate feedback for self-improvement.
+In addition, we expand the instruction-tune dataset for SFT and adopt a curriculum-based training approach, that is particularly effective in the gradual learning of complex video-text relationships.
+In our empirical validations, the VLM-RLAIF significantly outperforms previous models in multiple multimodal video-text understanding benchmarks, which implies good generalization performance across tasks.
+Limitations
+Given that our approach utilizes feedback synthesized by the AI model, the effectiveness of our proposed VLM-RLAIF largely depends on the quality of the AI model’s generated responses.
+In light of recent studies exploring the use of artificially generated data Koo et al.
+(2023); Das et al.
+(2024), we believe that there needs further research to enhance the quality of synthetically generated data, thereby establishing a more reliable RLAIF system.
+In addition, although we have evaluate our model across a range of benchmarks for VLMMs, e.g., videoQA, video-based generative tasks, retrieval, and recognition tasks, there are other tasks such as temporal reasoning Liang et al.
+(2022) that are necessary for VLMMs to be effectively applied in real-world scenarios.
+Application of our method to these tasks would be a great future research avenue.
+Acknowledgment
+This work was partly supported by the NRF grant (No.2022R1A2C4002300, 20%) and IITP grants (No.RS-2022-II220077 (10%), No.RS-2022-II220113 (25%), No.RS-2022-II220959 (10%), No.RS-2022-II220871 (10%), No.RS-2022-II220113 (10%, Yonsei AI), No.RS-2021-II211343 (5%, SNU AI), No.RS-2020-II201361 (5%, AI Innov.
+Hub), No.RS-2022-11220951 (5%)) funded by the Korea government (MSIT), NCSOFT and Artificial intelligence industrial convergence cluster development project funded by MSIT and Gwangju Metropolitan City.
+Appendix
+Appendix A Details About Object-Centric Instruction-Tune Data
 To enhance the alignment of Large Language Models (LLMs) with video content, we prioritize acquiring more video-text instruction-tune dataset that captures both the visual and temporal complexities of videos.
@@
 An example of the generated instruction-tune data is depicted in Fig.7-(b), showcasing the approach’s effectiveness in creating contextually rich instructional content.
-Appendix C Input Prompts for Preference Modeling
-We leverage three types of input prompts for context-aware preference modeling, as discussed in Sec. 3.2. Figures 89 and 10 show the detailed prompt example for each stage, respectively.
-Appendix D Various LLM Size for Reward Model
-In Tables 7 and 8, we assess the performance of VLM-RLAIF using varying reward model(RM) sizes, each initialized from VLM-RLAIF-7B and VLM-RLAIF-13B.
-In both settings, we use 7B VLM-SFT to initialize the policy model.
-For both sizes of RM, our RLAIF method showed a performance gain with a large margin from VLM-SFT across all evaluation methods.
-Specifically, RLAIF with 7B RM acquired 5-12% improvement on zero-shot video question answering.
-Scaling up the RM from 7B to 13B result in more improved performance, except for text-to-video retrieval task R@5 metric on MSR-VTT.
+Appendix B Input Prompts for Reward Modeling
+We leverage three types of input prompts for context-aware reward modeling, as discussed in Sec.
+3.1.
+Figure 8 presents a designed input prompt given to our method, resulting in two responses.
+Furthermore, Figure 9 illustrates the detailed input prompt used to select a preference between two responses.
+Additionally, Figure 10 displays the input prompt employed for generating a reward score.
 Figure 7: Designed input prompt for object-centric instruction-tune dataset generation from video narrative dataset and the data sample.
-In (a), the prompt we used to generate instruciton-following dataset is displayed in a white box.
+In (a), the prompt we used to generate instruction-following dataset is displayed in a white box.
 It includes the system prompt, a task definition, guidelines, principles, and an example set of input and desired output.
@@
 The VLM-RLAIF yields responses that are more accurately grounded in the visual input, as highlighted in blue for accurate responses and red for less accurate ones from VLM-SFT.
+Figure 12: Predictions of SFT model and the predicted reward.
+We present qualitative examples of questions and two responses generated from the VLM-SFT, and the reward given to each answer using the trained RM.
+The reward is higher for the answer with more accurately grounded answers, as highlighted in blue for accurate responses and red for less accurate contents.